Circular No. 7623

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 J1 S. Pravdo, E. Helin, and K. Lawrence, Jet Propulsion Laboratory,

report the discovery of a comet by NEAT on CCD images taken with the 1.2-m reflector at Haleakala. The object appears diffuse also on confirming CCD observations taken by L. Šarounová at Ondřejov (coma diameter ∼ 15′′) and by M. Tichý and J. Tichá at Kleť (coma diameter ∼ 8′′–10′′).

2001 UT α_{2000} δ_{2000} m_{1} Observer
May 11.25299 8^{h}38^{m}28.^{s}54 +13^{o}37′10′′.7 19.9 NEAT
11.83708 8 41 23.41 +13 31 55.0 17.3 Šarounová
11.85397 8 41 28.03 +13 31 49.7 17.1 Tichý

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD B. Sugerman, S. Lawrence, and A. Crotts, Columbia University, report

on evolution and new structure in the ejecta’s interaction with the inner equatorial ring (ER) of the remnant of SN 1987A: “Hubble Space Telescope (HST) STIS F28X50LP imaging and G750M spectroscopy (52′′ × 2′′ slit) were obtained on Apr. 27 and compared to earlier STIS observations. Using the notation of Lawrence et al. (2000, Ap.J. 537, 123), all nine spots reported on IAUC 7520 are reconfirmed, and of the four marginal detections reported therein, three are clearly detectable in imaging and the Hα spectrum, which we report with designation, p.a., radius, and the F28x50LP flux (counts/s): HS 12-050, 50^{o}, 0′′.63, 13.0; HS 13-165, 165^{o}, 0′′.52, 2.3; HS 14-249, 249^{o}, 0′′.72, 9.2. The fourth spot reported on IAUC 7520 remains marginally detected in the Hα spectrum at p.a. ∼ 310^{o}. HS 1-029, HS 10-040, and HS 12-050 are strongly blended; however, crowded-field photometric techniques indicate that HS 1-029 has brightened by < 10 percent since Nov. 2000, revealing a distinct break in its light curve. The brightening rate of HS 2-104, HS 3-126, and HS 4-091 may be decelerating, as well. HS 11-355 and 12-050, first identified in Nov. 2000, have each brightened more in six months than HS 4-091 has since Jan. 1999, and HS 2-104 is now nearly half as bright as HS 1-029, demonstrating that the spots do not evolve uniformly. The remnant of SN 1987A has rapidly made the transition from a few, restricted hot spots to having many interaction sites distributed around the full circumference of the ER; many of these are unresolved and blended at the resolution of HST. Frequent monitoring is strongly recommended, with observations tailored to sample the entire interaction region.”

2001 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green